What is the pH of a solution that is prepared by adding 38.34 mL of 1.83 M HCl to 0.772 L of 0.315 M trimethylamine (a weak base)?


1
Expert's answer
2021-10-31T00:04:25-0400

"H\nC\nl\n(\na\nq\n)\n+\nB\n(\na\nq\n)\n\u2192\nC\nl\n^\u2212\n(\na\nq\n)\n+\nH\nB\n^+\n(\na\nq\n)"

Molar mass of HCl =36.458g/mol

1.8×36.458= 65.62g

"\frac{38.34
 
m
L
 
H
C
l}{65.2}

= 0.59mol"

Molar mass of trimethylamine = 59.11

0.315×59.11= 18.62

"782\/" 18.62= 42moles

Ka=5.9×10-6


-log 5.9×10-6=5.3

Log42/0.59=1.1

pH= 6.4